# Efficient Spectral Element Method for the Euler Equations on Unbounded Domains in Multiple Dimensions

### v1

@@ 4 Numerical Results
where $z_{s}=15000$ m, $\Delta\gamma=0.5\mbox{s}^{-1}$ and $z_{max}$ is the top of the absorbing layer.
Figure 11: Time-converged numerical horizontal velocity perturbation $u$ (left) and b) vertical velocity $w$ (right) for the Linear Hydrostatic Mountain. The numerical solutions are displayed at t=30,000 s for $\Delta x=500~{}$m, $\Delta z=178~{}$m, and $\Delta z_{\rm Laguerre}=923~{}$m. The contours of the horizontal velocity (left) are in the range -0.025 to +0.025 m/s an 0.005 m/s intervals. Those of the vertical velocity (right) range between -0.005 and +0.005 with interval 0.0005. The dashed line shows where $\Omega^{F}$ meets $\Omega^{S}$. In $\Omega^{S}$ there is only one element of order 14 along the vertical direction (see Fig. 1 for reference.)
Figure 12: Comparison of vertical velocity contours of the numerical solution (blue) of the linear hydrostatic mountain wave problem with an approximate linear Fourier solution (red). The contours are the same contours used in the vertical velocity plot of figure 11. The numerical solution compares favorably to the analytical approximation and displays the expected vertical wavelength $\lambda_{z}=2\pi U/N\sim 6.4$ km.
Figure 11 shows the contours of horizontal and vertical velocity of the time-converged numerical solution at $t=30,000~{}$s. Our solution is stable, the outgoing waves are effectively damped within $\Omega^{S}$, and our solution is physically meaningful when compared to other numerical solutions of atmospheric models using spectral elements see e.g. [39].
In order to complete the validation of the hydrostatic mountain wave simulation, we constructed a linear Fourier solution using the approach outlined in [38]. The vertical velocity is expressed as a Fourier integral under the Boussinesq approximation that is then evaluated using adaptive Gauss quadrature in wavenumber space. Since this solution neglects vertical variations in density, it is only valid for heights less than a scale height ($\sim$9 km). Figure 12 compares vertical velocity profiles of the numerically obtained solution (in blue) with the linear Fourier solution (in red). The figure presents shows a good overlap of the two solutions and the existing deviations are expected, given the deviations between the two models due to the analytical solution relying on the linearization of the euler equations and the Boussinesq approximation. Furthermore, this deviation is comparable to other atmospheric models.
Table 4 shows the time per time step for different configurations of the linear hydrostatic mountain test case. A fixed domain end point is maintained for the cases where no semi-infinite element is used and instead only the number of vertical elements is changed. Even while only using ten additional order 4 spectral elements in the vertical direction to replace the Laguerre semi-infinite element the time per time step remains about 10% higher than running a single order 20 Laguerre element. If we would seek to run the entire vertical domain at the same resolution as the finite domain of our simulations utilizing a Laguerre semi-infinite element, the cost increase would be nearly 50% higher. Considerably few additional vertical spectral elements (we estimate no more than three) would have to be used to obtain a similar time per time step as using a single semi-infinite element in this case and this would also yield a much lower order discretization within the absorbing layer.
@@
Table 4: Timings of linear hydrostatic mountain simulations with and without Laguerre semi-infinite elements in the absorbing layer. Simulations using the extended finite domain have the same physical domain size as simulations using semi-infinite elements, but the space that would be otherwise occupied by these elements is instead occupied by elements of the finite domain. In this case the number of elements in the $x$-direction $N_{x}$ remains the same but the number of elements in the $z$-direction $N_{z}$ is adjusted. The additional elements are all of order $N=4$ but the vertical resolution $\Delta z$ is allowed to change.
@@ 5.1 Mass conservation
As discussed in [41], mass conservation is a key ingredient to ensure accurate forecasts, such as correctly diagnosing surface pressure. A gradual, secular loss of mass may distort the resolved dynamics over long term time-integration. In addition, accurate surface pressure prediction depends on a proper mass budget [41]. Also, mass conservation is a prerequisite for conserving other invariants, such as momentum and total energy. Many spectral element atmospheric models, including HOMME [40] and NUMA [22], have been designed explicitly to conserve mass.
@@ 6 Conclusion
We presented a nodal continuous Galerkin approach for approximating problems on unbounded domains in two dimensions. Utilizing a tensor product Legendre-Laguerre basis, we enable a seamless transition between finite and semi-infinite domains. We demonstrate that this transition can be achieved by simply replacing one of the bases in a standard tensor product spectral elements approach with a scaled Laguerre function (SLF) basis using collocated Legendre-Gauss-Radau (LGR) interpolation/quadrature points . Consequently, a continuous Galerkin spectral element in the finite domain is coupled through the mDSS operator to an unbounded domain. In the unbounded domain, a standard Legendre basis on LGL points is used in the finite direction, and a SLF basis with LGR points is used in semi-infinite direction. This approach is employed to construct absorbing layers on the semi-infinite domain $\Omega^{S}$, showcasing its effectiveness in damping outgoing waves in one dimension with minimal reflection. Subsequently, we demonstrate the ability of this approach to achieve accurate solutions for the advection-diffusion and Helmholtz equations in two dimensions. Finally, we illustrate that this method can effectively build absorbing layers for the non-linear compressible Euler equations in two dimensions. We showcase its capability to damp outgoing gravity waves in the classical linear hydrostatic mountain case. Furthermore, we provide evidence that our approach is relatively inexpensive compared to more classical Rayleigh damping approaches for a variety of 1D and 2D cases, and this efficiency increases with the desired size of the unbounded domain.
The work presented here emphasizes the potential of utilizing Laguerre-based absorbing layers for simulating atmospheric flow and numerical weather prediction. It is crucial to highlight the substantial time savings, particularly in large high-resolution simulations. The time typically allocated for computations in the absorbing layer could instead be directed towards the domain of interest or even discarded, favoring quicker simulations. This study does not aim to establish criteria for selecting the appropriate scaling parameter; hence, we limit ourselves to merely altering the domain size for test cases comparable with the literature. In this regard, the work presented in [46, 45, 6] opens up the possibility of time-evolving adaptive unbounded domains for simulating atmospheric flows. Additionally, we aspire to extend this model to 3D and employ it for efficient moist atmosphere simulations [42]. Lastly, we look forward to expanding this method to cases requiring an absorbing layer in both the lateral and vertical dimensions, and exploring Laguerre-Laguerre semi-infinite elements.
@@ 7 Code availability
The open source code Jexpresso used for this paper is available on Github [30]. All of the tests described in this paper can be accessed and run from the code at this link:https://github.com/smarras79/Jexpresso/tree/yt/Laguerre_test_suite. Timings were obtained by running the code swith Julia 1.9.3 on a Macbook Air M1 2020, with macOS Big Sur Version 11.6.
@@ 8 Acknowledgments
Tissaoui and Marras gratefully acknowledge the partial support by the National Science Foundation through NSF grant PD-2121367. Kelly gratefully acknowledges the support of the Office of Naval Research under grant # N0001419WX00721.
Appendix
Constructing the element right hand side

### v3

@@ 4 Numerical Results
where $z_{s}=15000$ m, $\Delta\gamma=0.1\mbox{s}^{-1}$ and $z_{max}$ is the top of the absorbing layer. The lateral boundaries are periodic and lateral damping layers are used on each side (these damping layers are in $\Omega^{F}$ and are not on semi-infinite elements). These boundary conditions are also used for the remaining other mountain test presented in this paper.
Figure 11 shows the contours of the vertical velocity at $t=30,000~{}$s. The solution is stable, the outgoing waves are effectively damped within $\Omega^{S}$, and the solution is physically meaningful when compared to other numerical solutions of atmospheric models using spectral elements see e.g. [46].
Figure 11: Time-converged numerical vertical velocity $w$ for the Linear Hydrostatic Mountain. The numerical solutions are displayed at t=30,000 s for $\Delta x=500~{}$m, $\Delta z=178~{}$m. The contours of the vertical velocity range between -0.005 and +0.005 with interval 0.0005. The dashed line shows where $\Omega^{F}$ meets $\Omega^{S}$. In $\Omega^{S}$ there is only one element of order 14 along the vertical direction (see Fig. 1 for reference.)
Table 4 shows the time per time step for different configurations of the linear hydrostatic mountain test case. A fixed domain end point $Z_{end}=30,000~{}$m is maintained for the cases where no semi-infinite element is used and instead only the number of vertical elements is changed. Even while only using ten additional order 4 spectral elements in the vertical direction to replace the Laguerre semi-infinite element the time per time step remains about 10% higher than running a single order 20 Laguerre element. If we would seek to run the entire vertical domain at the same resolution as the finite domain of these simulations utilizing a Laguerre semi-infinite element, the cost increase would be nearly 50% higher. Considerably few additional vertical spectral elements (we estimate no more than three) would have to be used to obtain a similar time per time step as using a single semi-infinite element in this case and this would also yield a much lower order discretization within the absorbing layer.
@@
Table 4: Timings of linear hydrostatic mountain simulations with and without Laguerre semi-infinite elements in the absorbing layer. A fixed vertical domain end point is used for simulation using only a finite domain while the simulations using semi-infinite elements are allowed to have varying $Z_{end}$ depending on the order of the semi-infinite elements. In this case the number of elements in the $x$-direction $N_{x}$ remains the same but the number of elements in the $z$-direction $N_{z}$ is adjusted. The additional elements are all of order $N=4$ but the vertical resolution $\Delta z$ is allowed to change.
In order to complete the validation of the hydrostatic and mountain wave simulations, we constructed a linear Fourier solution using the approach outlined in [45]. The vertical velocity is expressed as a Fourier integral under the Boussinesq approximation that is then evaluated using adaptive Gauss quadrature in wavenumber space. Since this solution neglects vertical variations in density, it is only valid for heights less than a scale height ($\sim$9 km). Figure 12 compares vertical velocity profiles of the numerically obtained solution (in blue) with the linear Fourier solution (in red). The figure shows a good overlap of the two solutions and the existing deviations are expected. This is given that the two models deviate due to the analytical solution relying on the linearization of the Euler equations and the Boussinesq approximation. Furthermore, this deviation is comparable to other atmospheric models.
Figure 12: Comparison of vertical velocity contours of the numerical solution (blue) of the linear hydrostatic mountain wave problem with an approximate linear Fourier solution (red). The contours are the same contours used in the vertical velocity plot of figure 11. The numerical solutions compare favorably to the analytical approximation and display the expected vertical wavelength $\lambda_{z}=2\pi U/\mathcal{N}\sim 6.4$ km. The Fourier solution is found under specific assumptions that are not met by the numerical model. For this reason, the small differences between the two contour sets are expected.
Schär mountain waves
The Schär test [42] consists of a uniform flow with a reference horizontal velocity $U=10~{}\rm{ms^{-1}}$ in a stratified atmosphere with Brunt-Väisälä frequency $\mathcal{N}=0.01$ $\rm{s^{-1}}$, sea-level pressure $p_{0}=1000~{}$hPa and and potential temperature at sea-level $\theta_{0}=280~{}$K. The flow impinges the five-peak mountain defined as
$z=he^{-\left(\frac{x}{a}\right)^{2}}\cos^{2}\left(\frac{\pi x}{\lambda_{c}}\right)$
with parameters $h=250$ m, $a=5000$ m, and $\lambda_{c}=4000$ m.
We consider a finite domain $\Omega^{F}=[-25,25]~{}{\rm km}\times[0,15]~{}$km, and discretize it using $N_{x}=20$ elements in the x direction and $N_{z}=7$ elements in the z direction. These elements are of polynomial order 10 in both directions, leading to an effective resolution $(\Delta x,\Delta z)=(214~{}{\rm m},250~{}{\rm m})$. A set of 20 Laguerre-Legendre semi-infinite elements are added on top of $\Omega^{F}$ which translates to the semi-infinite domain $\Omega^{S}=[-25,25]~{}{\rm km}\times[15,\infty)~{}$km. In the horizontal direction, each semi-infinite element uses an order 10 spectral element discretization using LGL nodes, and in the vertical direction it uses an order 14 Laguerre basis function on LGR nodes and a scaling factor $\lambda=300~{}$m which yields an end point $Z_{end}=28853$ m. Similarly to the previous mountain wave tests a Boyd-Vandeven filter [5] is used to help insure the stability of the simulations and the a sine squared function is used in the damping coefficient of the Rayleigh damping layer.
Figure 13: Schär mountain waves. Time-converged vertical velocity at t=36,000 s for $\Delta x=250~{}$m, $\Delta z=220~{}$m. The contours of the vertical velocity range between $-2.0~{}\rm{ms^{-1}}$ and $+2.0~{}\rm{ms^{-1}}$ with interval $0.1~{}\rm{ms^{-1}}.$ The dashed line shows where $\Omega^{F}$ meets $\Omega^{S}$. In $\Omega^{S}$ there is only one element of order 14 along the vertical direction (see Fig. 1 for reference.)
Figure 14: Comparison of vertical velocity contours of the numerical solution (blue) and semi-analytical Fourier solution (red) for the Schär mountain test. As for those shown in Fig. 12, some discrepancy is due to the linearization of the Euler equations used to calculate the semi-analytic Fourier solution.
Absorbing layer type $T^{*}$ $Z_{end}$ $T^{*}_{Finite}$ $T^{*}_{Laguerre}$ $N_{z}$
Semi-infinite elements of order 14 1 28853 88.62 11.38 7
Semi-infinite elements of order 18 1.03 33279 86.02 13.98 7
Extended finite domain of order 10 0.92 21000 100 N/A 8
1.03 9
1.14 10
1.27 11
Table 5: Timings of Schär mountain wave simulations with and without Laguerre semi-infinite elements in the absorbing layer. A fixed vertical domain end point is used for simulation using only a finite domain while the simulations using semi-infinite elements are allowed to have varying $Z_{end}$ depending on the order of the semi-infinite elements. In this case the number of elements in the $x$-direction $N_{x}$ remains the same but the number of elements in the $z$-direction $N_{z}$ is adjusted. The additional elements are all of order $N=10$ but the vertical resolution $\Delta z$ is allowed to change.
Figure 13 shows the vertical velocity of the time converged numerical solution at $t=36,000~{}$s. The solution is stable and outgoing waves are effectively damped withing $\Omega^{S}$. The solution is also physically meaningful and comparable to other numerical solutions of atmospheric models, see e.g [42]. Figure 14 shows an overlap of the numerical vertical velocity with a linear Fourier solution under the anelastic approximation [33, Eq.(A10)]. The figure shows a good overlap of the two solution and the existing deviations are within expectations given the differences between the models. Furthermore, these deviations are comparable to other atmospheric models.
Table 5 shows the time per time step for different configurations of the schär mountain problem. A fixed domain end point $Z_{end}=21,000~{}$m is maintained for the cases where no semi-infinite element is used and instead only the number of vertical elements is changed. For this case, even while using only 4 additional elements to replace the Laguerre semi-infinite element and resulting in nearly equivalent resolutions within the finite domain, the cost per time step is nearly $25\%$ higher than using semi-infinite elements of order 18 while also having a smaller domain height. Only two additional elements of order 10 can be used to replace the semi-infinite elements if a similar time to solution is desired, but this will come at the cost of the order of accuracy within the damping layer.
Remark: For the two mountain wave test cases, we have compared solutions obtained using a standard spectral element approach against solutions obtained using the proposed approach of using semi-infinite elements in the absorbing layer. We report the root mean squared error resulting from these comparisons in Appendix C and conclude that the proposed semi-infinite element approach does not introduce any significant errors relative to a standard spectral element approach.
@@ 5.1 Mass conservation
As discussed in [49], mass conservation is a key ingredient to ensure accurate forecasts, such as correctly diagnosing surface pressure. A gradual, secular loss of mass may distort the resolved dynamics over long term time-integration. In addition, accurate surface pressure prediction depends on a proper mass budget [49]. Also, mass conservation is a prerequisite for conserving other invariants, such as momentum and total energy. Many spectral element atmospheric models, including HOMME [47] and NUMA [27], have been designed explicitly to conserve mass. Table 6 presents the relative mass loss for the method with respect to the compressible Euler equations and its sensitivity to different time integration schemes. These tests were done on a rising-thermal bubble and demonstrate that the method can indeed conserve mass.
Time integrator Mass loss
SSPRK54 1.9545155453050947e-16
SSPRK33 1.1727093271830568e-15
MSRK5 3.9090310906101895e-16
Table 6: Relative mass loss for the compressible Euler equations and sensitivity to the time integrators of DifferentialEquations.jl for a rising-thermal bubble test. Mass conservation measured at t = 1000 s with active viscosity. Results are for inexact integration.
@@ 5.2 High-Altitude simulations
5.3 Dimensions and complexity
Since efficiency and simulation speed is a significant concern when performing numerical simulations, we analyze the complexity of the method for up to three dimensions.
The computational cost of spectral element methods may be reduced by 1) using a tensor-product basis and 2) evaluating derivative operations using sum factorization [10, Chapter 4]. Using these two strategies, the cost of evaluating a derivative in a $d$-dimensional domain along direction $1\leq i\leq d$ is $\mathcal{O}\left(N_{e}p_{1}p_{2}\cdots p_{i}^{2}\cdots p_{d}\right)$. In the standard SEM, where the same basis functions are used along each direction, this reduces to $\mathcal{O}\left(N_{e}p^{d+1}\right)$. Hence, the complexity increases by a factor of $p$ going from $d=2$ to $d=3$. This is one of the major reasons why most spectral element codes use a moderate basis function order $3\leq p\leq 6$ (see HOMME-NH [48], NUMA [17], etc.).
In 2D, let the semi-infinite domain $\Omega^{S}$ consist of $N_{x}$ semi-infinite elements with order $N_{LGL}$ in the finite direction and $N_{LGR}$ in the semi-infinite direction.
  1. 1.
    Cost of constructing a vertical derivative: $\mathcal{O}\left(N_{x}p_{LGL}p_{LGR}^{2}\right)$.
  2. 2.
    Cost of constructing a horizontal derivative: $\mathcal{O}\left(N_{x}p_{LGL}^{2}p_{LGR}\right)$.
In 3D, let the semi-infinite domain $\Omega^{S}$ consist of $N_{x}N_{y}$ semi-infinite elements.
  1. 1.
    Cost of constructing a vertical derivative: $\mathcal{O}\left(N_{x}N_{y}p_{LGL}^{2}p_{LGR}^{2}\right)$.
  2. 2.
    Cost of constructing a horizontal derivative: $\mathcal{O}\left(N_{x}N_{y}p_{LGL}^{3}p_{LGR}\right)$.
Going from 2D to 3D, the cost of constructing either a vertical or horizontal derivative asymptotically increases by a factor of $N_{y}p_{LGL}$. This is the same factor incurred in the finite domain $\Omega^{F}$ going from 2D to 3D. Hence, the semi-infinite element approach has a similar scaling with respect to the number of dimensions as the standard tensor product SEM utilizing Lagrange polynomials defined at LGL points.
@@ 6 Conclusion
We presented a nodal continuous Galerkin (CG) approach for discretizing PDEs on unbounded domains in one or two dimensions. Utilizing a tensor product Legendre-Laguerre basis, waves/disturbances may propagate between finite and semi-infinite domains without artificial reflections. This coupling is achieved by replacing one of the basis functions in a standard tensor product spectral elements approach with a scaled Laguerre function (SLF) using collocated Legendre-Gauss-Radau (LGR) interpolation/quadrature points. Consequently, a CG spectral element in the finite domain is coupled through the DSS operator to an unbounded domain. In the unbounded domain, a standard Legendre basis on LGL points is used in the finite direction, and a SLF basis with LGR points is used in semi-infinite direction. This approach is employed to construct absorbing layers on the semi-infinite domain $\Omega^{S}$, showcasing its effectiveness in damping outgoing waves in one dimension with minimal reflection. Subsequently, we demonstrate the ability of this approach to achieve accurate solutions for the advection-diffusion and Helmholtz equations in two dimensions. Finally, we illustrate that this method can effectively build absorbing layers for the non-linear compressible Euler equations in two dimensions. We showcase its capability to damp vertically propagating gravity waves in two standard idealized cases: a linear hydrostatic mountain and the Schär mountain. Furthermore, we demonstrate that the semi-infinite element approach is less expensive than the standard Rayleigh damping approach for a variety of 1D and 2D cases, and this efficiency increases with the desired size of the unbounded domain.
Reducing time-to-solution in large high-resolution simulations is crucial. The time typically allocated for computations in the absorbing layer could instead be directed towards the domain of interest or even discarded, favoring quicker simulations. This study does not aim to establish criteria for selecting the appropriate scaling parameter; hence, we limit ourselves to merely altering the domain size for test cases comparable with the literature. In this regard, the work presented in [55, 54, 6] opens up the possibility of time-evolving adaptive unbounded domains for simulating atmospheric flows. Additionally, we are extending this model to 3D for efficient moist atmosphere simulations [50] and expanding this method to cases requiring an absorbing layer in both the lateral and vertical dimensions.
@@ 7 Code availability
The open source code Jexpresso used for this paper is available on Github [36]. All of the tests described in this paper can be accessed and run from the code at this link:
https://doi.org/10.5281/zenodo.11490036. Timings were obtained by running the code with Julia 1.9.3 on a Macbook Air M1 2020, with macOS Big Sur Version 11.6.
@@ 8 Acknowledgments
The authors are grateful to the comments by two anonymous reviewers and by Tommaso Benacchio (Danish Meteorological Institute), who all improved the paper. Tissaoui and Marras gratefully acknowledge the partial support by the National Science Foundation through NSF grant PD-2121367. Kelly gratefully acknowledges the support of the Office of Naval Research under grant # N0001419WX00721. Makie [7] and VisIt [51] were used to create the plots.
Appendix A: Constructing the element right hand side
@@
Appendix B: Extension to 3D
The proposed semi-infinite element approach may be extended to 3D. Let $\mathbf{\xi}=(\xi,\eta,\zeta)$ be the coordinate of a point on the three dimensional reference element. Eq. (28) is generalized to three dimensions as follows:
$\psi_{l}(\mathbf{x})=h_{i}[\xi(\mathbf{x})]\otimes h_{j}[\eta(\mathbf{x})]\otimes\overline{h}_{k}[\eta(\mathbf{x})],\hskip 5.0pt\hskip 3.0ptl=i+(j-1)N_{LGL}+(k-1)N_{LGL}^{2},$ (49)
where $i\in\{1,\dots,N_{LGL}\}$, $j\in\{1,\dots,N_{LGL}\}$, and $k\in\{1,\dots,N_{LGR}\}$. It is then simple to extend equation (17) to three dimensions by adding an additional sum over the LGL nodes and including their corresponding weights in the product:
[…]
The remainder of the extension, such as the construction of the mass matrix can be done by following the same approach.
Appendix C: Error estimates for the linear hydrostatic mountain and Schär mountain cases
We present here the root mean squared error (RMSE) obtained from comparing two simulations of the linear hydrostatic mountain (LHM). The first is a standard spectral element simulation without any semi-infinite elements and the second uses semi-infinite elements of order 14 in the absorbing layer. The resolutions are the same for the finite domains in both simulations which use spectral elements of order 4. The domain extents are also the same for both simulations. The same comparison is also performed for the Schär mountain test case but with the order of elements within the finite domain being 10.
The RMSEs for both cases and all the variables are presented in table 7 for the solution at $t=10$ hours.
Variable $\rho$ $u$ $v$ $\theta$
RMSE (LHM) 2.50e-7 3.89e-4 3.71e-5 3.18e-4
RMSE (Schär) 5.95e-6 1.54e-2 3.50e-3 4.80e-3
Table 7: Root mean squared error for each variable at $t=10~{}$hours. The error is obtained by comparing a simulation using a standard spectral element approach against a simulation using semi-infinite elements of order 14 within the absorbing layer.
For both mountain wave cases, the order of magnitude of the errors shown in table 7 are similar or smaller than those obtained by Giraldo and Restelli in 2008 ([18]) who compare their spectral element method against the linearized analytical solutions. This makes us confident that the proposed approach does not introduce any significant errors compared to a standard spectral element approach for these two test cases. Furthermore, for LHM case, the total time to solution for the standard spectral element approach was $\sim 39,000~{}$ seconds, while the approach utilizing semi-infinite elements had a total time to solution of $\sim 24,000~{}$ and is thus $\sim 1.6$ times faster. For the Schär mountain case, the total time to solution for the standard spectral element approach was $\sim 21,000~{}$ seconds, while the approach utilizing semi-infinite elements had a time to solution of $\sim 13,000~{}$ seconds and is thus also $\sim 1.6$ times faster.